OpenStudy (anonymous):

What are the solutions of the system? y=x^2-3x+2 y=4x-4

OpenStudy (anonymous):

ok... do you want me to walk through this with you?

OpenStudy (anonymous):

set the equations equal to each other: x^2-3x+2 = 4x-4 make the right side equal to zero by subtractin 4x and adding 4 to both sides: x^2-3x+2 - 4x + 4 = 4x-4 - 4x + 4 x^2 - 7x + 6 = 0 are you ok up to this point?

OpenStudy (anonymous):

my choices are asking for the points u would graph and one says no solution.